# Differentially Private Stream Processing at Scale

### v2

@@ 3.4 Hierarchical Perturbation
4 System Implementation and Optimization
Figure 5: Execution of streaming differential privacy mechanism
@@ 4.1 State Management
In DP-SQLP, the system state store is a persistent storage system, backed by Spanner database [13] that provides high availability and fault tolerance. The system state store is co-managed by the DP-SQLP operator and the streaming framework for state update and maintenance. All the state information required by streaming differential privacy is stored in the system state store. For a pictorial depiction, see Figure 2.
@@
  2. 2.
    Cross-user aggregation: Records in one micro-batch are grouped by key and aggregated, which form a delta result [23]. After that, the delta result will be merged into the data buffer that is loaded from the system state store.
  3. 3.
@@
4.2 Parallel Execution
Figure 6: Parallel execution within DP-SQLP operator
When building the DP-SQLP operator, we leverage the F1 query engine [40] for its wide range of data sources and distributed query execution (Figure 6). The user contribution bounding step is executed by the user contribution bounding server. The privacy key selection and hierarchical perturbation are executed by the data perturbation server. Both servers contain thousands of workers that are horizontally scalable. Input data is first read by F1 query engine, partitioned, then sent to the user contribution bounding server through Remote Procedure Calls (RPCs). After that, the bounded data will stream back to F1, re-partitioned, then being sent to data perturbation server for key selection and hierarchical perturbation.
4.3 Empty Key Release Prediction
Since the data stream is unordered and unbounded, the existence of user contributions within each micro-batch can be arbitrary, as shown in Figure 7. It is possible that some keys do not have any user records in a micro-batch. In the traditional streaming systems, states are not updated unless new records appear in the micro-batch[45]. Therefore, the system only needs to load states with keys from the current micro-batch.
@@ 4.3 Empty Key Release Prediction
Figure 7: Some key may not have records within certain micro-batch
  1. 1.
@@
4.4 Privacy Accounting for DP-SQLP
@@
  • Privacy accounting for streaming key selection: When executing Algorithm 1 in DP-SQLP, the value added to each leaf node is the unique user count. Therefore, the per-user sensitivity for each DP-Tree is one. In addition, we restart the Algorithm 1 once a key is selected, and data accumulated in the data buffer is released immediately. As a result, each user may participate in at most $C$ rounds of key selection. Given the $(\varepsilon,\delta)$ privacy budget for each round of Algorithm 1, the total privacy cost for streaming key selection is calculated using the empirically tighter variant of advanced composition [19] with $C$-fold.
  • Privacy accounting for hierarchical perturbation: For each user, in the worst case, all $C$ contributions can go to the same node of a single DP-Tree , we scale up the sensitivity corresponding to any single node in the tree to $L_{1}=C\cdot L$, and ensure that each tree still ensures $\rho$-zCDP. After that, we use the conversion from [37, Proposition 3] to translate privacy cost from $\rho$-zCDP to $(\varepsilon,\delta)$-DP guarantee  [^8] The exact computation is from https://github.com/IBM/discrete-gaussian-differential-privacy/blob/master/cdp2adp.py#L123.
@@ 5 Experiments
For both baseline 1 and baseline 2, the one-shot differential privacy mechanism is executed by Plume [4] with the Gaussian mechanism. Each one-shot differential privacy execution guarantees $(\varepsilon,\delta)$-differential privacy. We also adopt the optimal composition theorem for DP [30] to maximize the baseline performance.
@@ 5.1 Synthetic Data
Table 1: Data utility measure with synthetic data $(\varepsilon=6,\delta=10^{-9})$
@@
5.2 Reddit Data
In the next step, we apply the same experiment to real-world data. Webis-tldr-17-corpus [43] is a popular dataset consisting of 3.8 million posts associated with 1.4 million users on the discussion website Reddit. Our task is to count the user participation per subreddit (specific interest group on Reddit).
@@ 5.2 Reddit Data
Table 2: Data utility measure with the Reddit data $(\varepsilon=6,\delta=10^{-9})$
@@
Figure 8: Micro-batch execution time $(\varepsilon=6,\delta=10^{-9})$
5.3 Execution Performance
@@
5.4 Parameter Tuning
Figure 9: Metrics under different contribution limit $(\varepsilon=6,\delta=10^{-9})$
@@ 5.4 Parameter Tuning
6 Industry Applications
6.1 Apply DP-SQLP to Google Shopping
@@
6.2 Streaming DP in Google Trends
@@
7 Conclusion and Future Work
In this paper, we presented a streaming differentially private system (DP-SQLP) that is designed to continuously release DP histograms using existing distributed stream processing systems. With user contribution bounding, streaming key selection, and hierarchical perturbation, we provide a formal $(\varepsilon,\delta)$-user level DP guarantee for arbitrary data streams. In addition to the algorithmic design, we implemented our system using a streaming framework similar to Spark streaming, Spanner database, and F1 query engine from Google. The experiments were conducted using both synthetic data and Reddit data. We compared DP-SQLP with two baselines that use one-shot differential privacy algorithms (and privacy accumulation was composed over time). The experiment results demonstrated a significant performance improvement in terms of data utility. In the end, we present two industry applications that apply DP-SQLP and the streaming private key selection algorithm to the production use cases.
There are three main ways in which our system can be further extended. First, due to the nature of the underlying streaming system (i.e., an input driven stream), we cannot protect the event times of a user. It would be interesting to redesign stream processing systems that are compatible with DP even in the context of protecting the event timestamps. Second, our algorithms are primarily designed to provide a centralized DP guarantee, where the final outcome of the system is guaranteed to be DP. It is worth exploring DP streaming system designs that allow stronger privacy guarantees like pan-privacy [18]. Third, we bound the contribution of each user globally by $C$. However, for higher fidelity, it is important to explore approaches to perform per-key contribution bounding. Naive approaches that address this issue can get complicated due to the fact that we are dealing with an input driven stream.
Acknowledgement
@@ Appendix A Terms from Streaming Data Processing
For a stream processing system, the statistic release is driven by inputs. The output computation is based on the current input (stateless), or a sequence of inputs (stateful) [45].
@@
Event time is the time at which the event itself actually occurred, i.e. a record of system clock time (for whatever system generated the event) at the time of occurrence. Each event timestamp is determined from a discrete, ordered domain, which is called event time domain [42].
@@
Windowing slices up a dataset into finite chunks for processing as a group. It determines where in event time data are grouped together for processing [3]. Some common windows include fixed windows, sliding windows and sessions.
@@
Trigger is a mechanism to stimulate the output of a specific window at a grouping operation. It determines when in processing time the results of groupings are emitted as panes [3]. There are many ways to define a trigger, including watermark, processing time timers and data arrival.
@@
Each micro-batch is a partitioned dataset. The continuous data stream is discretized into many micro-batches [3], and the execution of each micro-batch resembles continuous processing.
@@ Appendix B Differential Privacy on Streams
Definition B.2 (Differential Privacy [16, 15]).
@@
We can similarly define Concentrated DP (zCDP) [20, 7] with the neighbouring relation given by Definition B.1. We use zCDP as an analysis tool [^12] Throughout the paper, zCDP is used as a convenient tool to analyze DP-Binary Tree. The privacy guarantee is reported in $(\varepsilon,\delta)$-DP., as it has clean composition properties, but we convert the final guarantee into $(\varepsilon,\delta)$-DP using standard tools. We use three facts:
@@ Appendix C Privacy Under Continual Observation and Binary Tree Aggregation
Consider a data stream $D=x_{1},\ldots,x_{n}$ with each $x_{i}\in[0,L]$. The objective is to output $\left\{S_{j}=\sum\limits_{i=1}^{j}x_{i}\right\}$ while preserving DP (with the neighborhood relation defined w.r.t. changing any one of the $x_{i}$’s). We will use the so-called binary tree aggregation algorithm [17, 11, 26] stated below.
Algorithm 4 DP-Binary Tree Aggregation (exposition from [29])
@@
Figure 10: DP-Tree with nodes encoding
@@
The proof of Theorem C.2 follows immediately from the tail properties of Gaussian distribution. One can improve the constants in the guarantee by using variance reduction techniques from [26], which we briefly describe below.
Honaker estimation for variance reduction: The idea of Honaker estimation [26] is to use multiple estimates for the information of same node in the binary tree for variance reduction. Consider any node $\texttt{node}_{i}$, and the information that is inserted into it by Algorithm 4. Let $\mathcal{T}_{\texttt{node}_{i}}$ be the subtree rooted at node $\texttt{node}_{i}$. Notice that sum of all the nodes at each level of $\mathcal{T}_{\texttt{node}_{i}}$ is an unbiased and independent estimate of the information contained in $\texttt{node}_{i}$. Furthermore, we also know the variance of such an estimate. For example, if a level $\mathcal{T}_{\texttt{node}_{i}}$ has $m$ nodes, then the variance of the estimate from that level is $m\sigma^{2}$. We can use this information to reduce the variance in the estimate of $\texttt{node}_{i}$ The following is the formalization of this idea.
@@
${\sf Variance}({\sf node}_{i})=\left(\sum\limits_{j=0}^{\kappa-1}c_{j}^{2}\cdot 2^{j}\right)\cdot\sigma^{2}=\frac{1}{2\cdot(1-2^{-\kappa})}\cdot\sigma^{2}.$ (1)
@@
$\left(S^{\sf priv}_{i}-S_{i}\right)\sim\mathcal{N}\left(0,\sum\limits_{j=1}^{m}{\sf Variance}({\sf node}_{j})\right).$ (2)
@@
Although existing works commonly assume that each user corresponds to a single record, in practice, a single user can contribute multiple records. Thus, we must account for this in the DP analysis. The simplest approach – which we follow – is to limit the number of contributions per person to some constant $C$ within the entire data stream $D$. If we have a DP guarantee for a single record, we can apply group privacy to obtain a DP guarantee for $C$ records. Specifically, if we have $(\varepsilon,\delta)$-DP with respect to the addition or removal of one record, then this implies $\left(C\cdot\varepsilon,\frac{e^{C\cdot\varepsilon}-1}{e^{\varepsilon}-1}\delta\right)$-DP with respect to the addition or removal of $C$ records [46]. The group privacy is used in privacy accounting for hierarchical perturbation (Section 4.4).
@@ Appendix D Missing Details from Algorithm 1
Setting $\sigma$ in Line 1: The binary tree has at most $T=|\texttt{Tr}|$ nodes. By Theorem C.1, the tree aggregation algorithm satisfies $\rho=\frac{\lceil\log_{2}(T)\rceil}{2\sigma^{2}}$-zCDP. Now, we translate the zCDP guarantee into $(\varepsilon,\delta)$-DP (with $\varepsilon$ being a parameter of $\sigma$) via [37, Proposition 2.3]  [^13] The exact computation is from https://github.com/IBM/discrete-gaussian-differential-privacy/blob/master/cdp2adp.py#L123. Using this we obtain $\sigma$ as a function of fixed $\varepsilon$ and $\delta$ and $T$, when used as parameter in Algorithm 1 is guaranteed to ensure $(\varepsilon,\delta)$-DP for the binary tree used there.
@@
$\hat{q}_{{tr_{i},k}}-\texttt{count}_{k}(D_{tr_{i}})\sim\mathcal{N}\left(0,\underbrace{\sum\limits_{j=1}^{k}{\sf Variance}({\sf node}_{j})}_{\lambda^{2}}\right).$ (3)
@@
$\begin{split}{\underset{\mathcal{T}_{k}}{\mathbb{P}}\left[\forall tr_{i}\in\texttt{Tr}~{}~{}|\hat{q}_{{tr_{i},k}}-\texttt{count}_{k}(D_{tr_{i}})|\leq\sqrt{\frac{2\ln(T/\beta)\lceil\lg(T)\rceil\sigma^{2}}{\pi}}\right]}\\ \geq 1-\beta.\end{split}$ (4)

### v3

@@ 3.4. Hierarchical Perturbation
4. System Implementation and Optimization
Figure 5. Execution of streaming differential privacy mechanism
@@ 4.1. State Management
In DP-SQLP, the system state store is a persistent storage system, backed by Spanner database (Corbett et al., 2013) that provides high availability and fault tolerance. The system state store is co-managed by the DP-SQLP operator and the streaming framework for state update and maintenance. All the state information required by streaming differential privacy is stored in the system state store. For a pictorial depiction, see Figure 2.
@@
  2. (2)
    Cross-user aggregation: Records in one micro-batch are grouped by key and aggregated, which form a delta result (Fegaras, 2016). After that, the delta result will be merged into the data buffer that is loaded from the system state store.
  3. (3)
@@
4.2. Parallel Execution
Figure 6. Parallel execution within DP-SQLP operator
When building the DP-SQLP operator, we leverage the F1 query engine (Samwel et al., 2018) for its wide range of data sources and distributed query execution (Figure 6). The user contribution bounding step is executed by the user contribution bounding server. The privacy key selection and hierarchical perturbation are executed by the data perturbation server. Both servers contain thousands of workers that are horizontally scalable. Input data is first read by F1 query engine, partitioned, then sent to the user contribution bounding server through Remote Procedure Calls (RPCs). After that, the bounded data will stream back to F1, re-partitioned, then being sent to data perturbation server for key selection and hierarchical perturbation.
4.3. Empty Key Release Prediction
Since the data stream is unordered and unbounded, the existence of user contributions within each micro-batch can be arbitrary, as shown in Figure 7. It is possible that some keys do not have any user records in a micro-batch. In the traditional streaming systems, states are not updated unless new records appear in the micro-batch(To et al., 2018). Therefore, the system only needs to load states with keys from the current micro-batch.
@@ 4.3. Empty Key Release Prediction
Figure 7. Some key may not have records within certain micro-batch
  1. (1)
@@
4.4. Privacy Accounting for DP-SQLP
@@
  • Privacy accounting for streaming key selection: When executing Algorithm 1 in DP-SQLP, the value added to each leaf node is the unique user count. Therefore, the per-user sensitivity for each DP-Tree is one. In addition, we restart the Algorithm 1 once a key is selected, and data accumulated in the data buffer is released immediately. As a result, each user may participate in at most $C$ rounds of key selection. Given the $(\varepsilon,\delta)$ privacy budget for each round of Algorithm 1, the total privacy cost for streaming key selection is calculated using the empirically tighter variant of advanced composition (Dwork and Roth, 2014) with $C$-fold.
[…]
@@ 5. Experiments
For both baseline 1 and baseline 2, the one-shot differential privacy mechanism is executed by Plume (Amin et al., 2022) with the Gaussian mechanism. Each one-shot differential privacy execution guarantees $(\varepsilon,\delta)$-differential privacy. We also adopt the optimal composition theorem for DP (Kairouz et al., 2017) to maximize the baseline performance.
@@ 5.1. Synthetic Data
Table 1. Data utility measure with synthetic data $(\varepsilon=6,\delta=10^{-9})$
@@
5.2. Reddit Data
In the next step, we apply the same experiment to real-world data. Webis-tldr-17-corpus (Syed et al., 2017) is a popular dataset consisting of 3.8 million posts associated with 1.4 million users on the discussion website Reddit. Our task is to count the user participation per subreddit (specific interest group on Reddit).
@@ 5.2. Reddit Data
Table 2. Data utility measure with the Reddit data $(\varepsilon=6,\delta=10^{-9})$
@@
Figure 8. Micro-batch execution time $(\varepsilon=6,\delta=10^{-9})$
5.3. Execution Performance
@@
5.4. Parameter Tuning
@@
Figure 9. Metrics under different contribution limit $(\varepsilon=6,\delta=10^{-9})$
6. Industry Applications
6.1. Apply DP-SQLP to Google Shopping
@@
6.2. Streaming DP in Google Trends
@@
7. Conclusion and Future Work
In this paper, we presented a streaming differentially private system (DP-SQLP) that is designed to continuously release DP histograms. We provide a formal $(\varepsilon,\delta)$-user level DP guarantee for arbitrary data streams. In addition to the algorithmic design, we implemented our system using a streaming framework similar to Spark streaming, Spanner database, and F1 query engine from Google. The experiments were conducted using both synthetic data and Reddit data. We compared DP-SQLP with two baselines, and the results demonstrated a significant performance improvement in terms of data utility. In the end, we present two industry applications that apply DP-SQLP and the streaming private key selection algorithm to the production use cases.
There are three main ways in which our system can be further extended. First, in the design of the system we have used DP-tree aggregation (Chan et al., 2011; Dwork et al., 2010a; Honaker, 2015) as the baseline DP algorithm. In recent research (Henzinger et al., 2023; Denisov et al., 2022), it has been shown that DP-tree aggregation is (significantly) sub-optimal in terms of privacy/utility trade-off, compared to general matrix factorization based mechanisms (DP-MF) (Matousek et al., 2014). However, DP-MF algorithms are not in general compatible with systems operating on data streams. Recently, following to our work, (McMahan et al., 2024) provided a streaming variant of DP-MF. In future incarnations of our system, we plan to incorporate this.
Second, our algorithms are primarily designed to provide a centralized DP guarantee, where the final outcome of the system is guaranteed to be DP. It is worth exploring DP streaming system designs that allow stronger privacy guarantees like pan-privacy (Dwork et al., 2010b).
Third, we bound the contribution of each user globally by $C$. However, for higher fidelity, it is important to explore approaches to perform per-key contribution bounding. Naive approaches that address this issue can get complicated due to the fact that we are dealing with an input driven stream.
Acknowledgements.
@@ Appendix A Terms from Streaming Data Processing
For a stream processing system, the statistic release is driven by inputs. The output computation is based on the current input (stateless), or a sequence of inputs (stateful) (To et al., 2018).
@@
Event time is the time at which the event itself actually occurred, i.e. a record of system clock time (for whatever system generated the event) at the time of occurrence. Each event timestamp is determined from a discrete, ordered domain, which is called event time domain (Srivastava and Widom, 2004).
@@
Windowing slices up a dataset into finite chunks for processing as a group. It determines where in event time data are grouped together for processing (Akidau et al., 2015). Some common windows include fixed windows, sliding windows and sessions.
@@
Trigger is a mechanism to stimulate the output of a specific window at a grouping operation. It determines when in processing time the results of groupings are emitted as panes (Akidau et al., 2015). There are many ways to define a trigger, including watermark, processing time timers and data arrival.
@@
Each micro-batch is a partitioned dataset. The continuous data stream is discretized into many micro-batches (Akidau et al., 2015), and the execution of each micro-batch resembles continuous processing.
@@ Appendix B Differential Privacy on Streams
Definition B.2 (Differential Privacy (Dwork et al., 2006b; Dwork et al., 2006a)).
@@
[…]
@@ Appendix C Privacy Under Continual Observation and Binary Tree Aggregation
Consider a data stream $D=x_{1},\ldots,x_{n}$ with each $x_{i}\in[0,L]$. The objective is to output $\left\{S_{j}=\sum\limits_{i=1}^{j}x_{i}\right\}$ while preserving DP (with the neighborhood relation defined w.r.t. changing any one of the $x_{i}$’s). We will use the so-called binary tree aggregation algorithm (Dwork et al., 2010a; Chan et al., 2011; Honaker, 2015) stated below.
Algorithm 4 DP-Binary Tree Aggregation (exposition from (Kairouz et al., 2021))
@@
Figure 10. DP-Tree with nodes encoding
@@
The proof of Theorem C.2 follows immediately from the tail properties of Gaussian distribution. One can improve the constants in the guarantee by using variance reduction techniques from (Honaker, 2015), which we briefly describe below.
Honaker estimation for variance reduction: The idea of Honaker estimation (Honaker, 2015) is to use multiple estimates for the information of same node in the binary tree for variance reduction. Consider any node $\texttt{node}_{i}$, and the information that is inserted into it by Algorithm 4. Let $\mathcal{T}_{\texttt{node}_{i}}$ be the subtree rooted at node $\texttt{node}_{i}$. Notice that sum of all the nodes at each level of $\mathcal{T}_{\texttt{node}_{i}}$ is an unbiased and independent estimate of the information contained in $\texttt{node}_{i}$. Furthermore, we also know the variance of such an estimate. For example, if a level $\mathcal{T}_{\texttt{node}_{i}}$ has $m$ nodes, then the variance of the estimate from that level is $m\sigma^{2}$. We can use this information to reduce the variance in the estimate of $\texttt{node}_{i}$ The following is the formalization of this idea.
@@
(1) ${\sf Variance}({\sf node}_{i})=\left(\sum\limits_{j=0}^{\kappa-1}c_{j}^{2}\cdot 2^{j}\right)\cdot\sigma^{2}=\frac{1}{2\cdot(1-2^{-\kappa})}\cdot\sigma^{2}.$
@@
(2) $\left(S^{\sf priv}_{i}-S_{i}\right)\sim\mathcal{N}\left(0,\sum\limits_{j=1}^{m}{\sf Variance}({\sf node}_{j})\right).$
@@
Although existing works commonly assume that each user corresponds to a single record, in practice, a single user can contribute multiple records. Thus, we must account for this in the DP analysis. The simplest approach – which we follow – is to limit the number of contributions per person to some constant $C$ within the entire data stream $D$. If we have a DP guarantee for a single record, we can apply group privacy to obtain a DP guarantee for $C$ records. Specifically, if we have $(\varepsilon,\delta)$-DP with respect to the addition or removal of one record, then this implies $\left(C\cdot\varepsilon,\frac{e^{C\cdot\varepsilon}-1}{e^{\varepsilon}-1}\delta\right)$-DP with respect to the addition or removal of $C$ records (Vadhan, 2017). The group privacy is used in privacy accounting for hierarchical perturbation (Section 4.4).
@@ Appendix D Missing Details from Algorithm 1
[…]
@@
(3) $\hat{q}_{{tr_{i},k}}-\texttt{count}_{k}(D_{tr_{i}})\sim\mathcal{N}\left(0,\underbrace{\sum\limits_{j=1}^{k}{\sf Variance}({\sf node}_{j})}_{\lambda^{2}}\right).$
@@
(4) $\begin{split}{\underset{\mathcal{T}_{k}}{\mathbb{P}}\left[\forall tr_{i}\in\texttt{Tr}~{}~{}|\hat{q}_{{tr_{i},k}}-\texttt{count}_{k}(D_{tr_{i}})|\leq\sqrt{\frac{2\ln(T/\beta)\lceil\lg(T)\rceil\sigma^{2}}{\pi}}\right]}\\ \geq 1-\beta.\end{split}$